OpenStudy (anonymous):

Quadrilateral ABCD with vertices A(4, 3), B(4, -2), C(-4, -2) and D(-4, 3) is a rectangle, find the length of the diagonals.

OpenStudy (campbell_st):

use the distance formula A(4,3) and C(-4, -2) \[d=\sqrt{(x _{2}-x_{1})^2 + (y _{2} - y _{1})^2}\]

OpenStudy (anonymous):

would it be square root of 39 Im so cunfused

OpenStudy (campbell_st):

\[d = \sqrt{(4 - (-4))^2 + (3 - (-2))^2} = \sqrt{8^2 + 5^2}\]

OpenStudy (anonymous):

oh so its the square root of 89